Question HelpGardi Manufacturing uses normal costing for its​ job-costingsystem, which has two​ direct-cost categories​...

Gardi Manufacturing uses normal costing for its job-costing system, which has two direct-cost categories (direct materials and direct manufacturing labor) and one indirect-cost category (manufacturing overhead). The following information is obtained for 2017.

-Total manufacturing costs, $8,350,000

- Manufacturing overhead allocated, $4,100,000 (allocated at rate of $250% of direct manufacturing labor costs)

- Work in process inventory on January 1, 2017, $350,000

- Cost of finished goods manufactured, $7,949,000

1. Use information in the first to bullet points to calculate (a) direct manufacturing labor costs in 2017 and (b) cost of direct materials used in 2017.

2. Calculate the ending work in process inventory on December 31, 2017
Answer #1

1a) Direct manufacturing labor cost = 4100000/2.5 = 1640000

1b) Direct material used = 8350000-4100000-1640000 = 2610000

1c) Ending work in process = 350000+8350000-7949000 = 751000